OpenStudy (anonymous):

cos (50degrees) = 6.1/x solve for x

OpenStudy (anonymous):

Note that \(\cos(50^{\circ}) = \dfrac{6.1}{x} \implies x\cos(50^{\circ})=6.1 \implies x = \dfrac{6.1}{\cos(50^{\circ})}\approx \ldots\) I assume at this point you'll use a calculator to simplify this. Does this make sense? :-)

OpenStudy (ahsome):

\[cos(50)=\frac{6.1}{x}\]Multiply both sides by \(x\)\[x*cos(50)=6.1\]Divide both sides by \(cos(50)\)\[x=\frac{6.1}{cos(50)}\]\[x=\frac{6.1}{0.642}\]\[x=9.49\]

OpenStudy (anonymous):

ty both(:

OpenStudy (ahsome):

No problem @kah.x :)